# CPEG: Leveraging Consistency Policy with Consensus Guidance for Multi-agent Exploration

### v1

@@ V-C Time Efficiency
In addition to its superior performance, CPEG also possesses higher time efficiency compared with methods based on diffusion models. This efficiency is essential for practical scalability in multi-agent systems and for facilitating real-time decision-making processes. To demonstrate this, we replace the policy class in our method with a diffusion model (CPEG-DM) and an accelerated variant of it (CPEG-DPM) [33], and compared the time required by the different algorithms for each training iteration, as shown in Table II.
@@
In Navigation and HalfCheetah(2x3) scenarios, our method attains training speeds up to $\times6$ and 6.8×6.8$\times$6.8 × faster than those of diffusion-based methods, rivaling the time efficiency of previous methods like HATD3. Despite using DPM-Solver, CPEG-DPM still faces challenges in achieving satisfactory training efficiency. Moreover, the training speed of algorithms is greatly impacted by the scaling in MAS. We selected Humanoid(17x1) scenario, which involves 17 agents, for evaluating our methods under large-scale agents. In this scenario, the training time of diffusion-based methods significantly exceeds that of CPEG, demonstrating the scalability of our method in MARL.
@@
Due to the substantial reduction in sampling steps, the consistency model is inherently less expressive than the diffusion model [36]. To evaluate this, we conduct a performance comparison between CPEG and diffusion-based methods in three sparse scenarios. In comparison to CPEG-DM and CPEG-DPM, CPEG exhibits a performance decrease in MPE scenarios, yet outperforms it in the HalfCheetah(2x3) and Humanoid (17x1) environments. This phenomenon could potentially be attributed to that the consistency policy can effectively transmit action gradients in one step, thereby alleviating the challenges of policy training via Q-function in diffusion-based methods [34]. Moreover, the slight performance trade-off in CPEG is considered acceptable, given its significant improvements in training and online inference efficiency.
@@ V-D Qualitative Analysis and Visualization
Figure 5: Example results obtained by CPEG and baselines on exploration visitation. (a) shows a Reacher4 task, with four targets in different colors. (b)-(d) shows the state visitation of CPEG, HASAC and MAPPO. (e) counts the targets covered by policies during exploration.
To showcase the exploration capabilities of our method, consider a didactic example illustrated in Fig. 5(a), which demonstrates different MARL algorithms applied to the sparse Reacher4 task [23] featuring four targets. The agents consist of two-arm robots tasked with moving the robot’s end effector close to one of the four targets. The depth of color represents the magnitude of the state visitation probability. In this scenario, we configure sparse rewards, meaning that the agents receive rewards only upon making contact with all four targets. Besides, the agents need to cooperate in exploration to reach the target, and the dispersion of targets often leads to agents adopting one single policy, underscoring the necessity for multimodal policies and cooperative exploration. In Fig. 5(b), our method learns a multimodal joint policy, enabling the effector to reach four different targets relatively uniformly, without converging to a unimodal policy. Conversely, in Fig. 5(c) and (d), baselines primarily focus on visiting certain targets, thereby limiting exploration in the complex environment. Additionally, we quantify the targets covered by the policy during exploration in Fig. 5(e). The curves indicate that our method achieves a state visitation rate of 83%, compared to 62% for HASAC, reflecting a 20% improvement. These results demonstrate that CPEG can facilitate a more efficient exploration of the four different targets in the scenario, highlighting the importance of multimodal policies and cooperative exploration capabilities among agents.
Figure 6: Trajectory visualization in HalfCheetah (2x3). Different colors represent the corresponding consensus between agents.
In order to gain a better understanding of the consensus learner within CPEG, we generate visualizations of consensus guidance and observation embeddings. In Fig. 6, we visualize the true trajectories generated by the consensus-guided consistency policy in the HalfCheetah(2x3) scenario. The trajectories are color-coded to represent the corresponding consensus generated during execution. Across stages A, B, and C, although the observations of the two agents differ, they converge to a common consensus (the same color). This indicates that the consensus effectively reflects the global state, enabling multiple agents to coordinate their actions. By leveraging this shared consensus, multiple agents can cooperate more effectively, thereby enhancing their exploration of the environment. Additionally, we present the visualizations obtained by applying the PCA technique to the observation embeddings. The goal is to show whether the consensus learner can deduce distinguishable consensus from the local observations, aiding in the guidance of policy during the exploration process. As shown in Fig. 7, the visualizations demonstrate that the consensus learner can generate informative discrete representations by identifying potential correlations within individual observations of agents.
@@
Figure 7: Visualization of the observation embeddings. We choose the last $30k$ steps for visualization. The colors denote different consensuses.
@@ V-E Ablation Study
In this subsection, we conduct ablation studies in Navigation scenario to investigate the impact of three main components in CPEG: (1) replacing the consistency policy with a deterministic policy (similar to MADDPG), while incorporating consensus guidances and observations directly as the inputs (CPEG-w/o-CP), (2) removing the consensus learner module, forcing agents to rely solely on their own observations for exploration during training (unguided exploration) and achieving objectives during execution (CPEG-w/o-CG), and (3) removing the self-reference mechanism, which relaxes the constraint on consistency policies (CPEG-w/o-SR). The experimental results, as shown in Fig. 8, reveal that CPEG-w/o-CP quickly converges to a locally optimal policy. This result indicates that the consistency policy significantly enhances the exploration, preventing agents from converging to local optima. Additionally, CPEG outperforms CPEG-w/o-CG, which does not use consensus guidance, in both convergence rate and final performance. Similarly, when using only consensus guidance, CPEG-w/o-SR exhibits a lack of exploitation of past successes, reducing sample efficiency and leading to performance degradation. Overall, the results indicate that these elements collectively contribute to the improved performance of our method.
@@ VI Conclusions and Future Work
In this paper, we propose CPEG, a novel multi-agent exploration method with a consensus-guided consistency policy that facilitates cooperative exploration and a self-reference mechanism that constrains the generated actions. To the best of our knowledge, our study represents the first effort in leveraging consistency policies in MARL. The empirical results from six environments demonstrate that CPEG performs comparably to baseline algorithms in dense reward settings while in sparse reward settings, which pose more challenges for exploration, our algorithm outperforms baselines by 20%.
This work is the first to showcase the advantages of consensus-guided consistency policies in the realm of multi-agent exploration. On the other hand, the present work is confined to exploration within single-task scenarios. In the future, we intend to study how to better harness the multimodal nature of consistency policies to devise a faster and more efficient algorithm for multi-task MARL.

### v2

@@ V-C Time Efficiency
Alongside its superior performance, we observe that CPIG also possesses higher time efficiency compared with other methods based on diffusion models. This efficiency is essential for practical scalability in multi-agent systems and for facilitating real-time decision-making processes. To demonstrate this, we replace the policy class in our method with a diffusion model (CPIG-DM) and an accelerated variant of it (CPIG-DPM) [36], and compared the time required by the different algorithms for each training iteration, as shown in Table II.
@@
In Navigation and HalfCheetah(2x3) scenarios, our method attains training speeds up to $\times6$ and 6.8×6.8$\times$6.8 × faster than those of diffusion-based methods, rivaling the time efficiency of previous methods like HATD3. Despite using DPM-Solver, CPIG-DPM still faces challenges in achieving satisfactory training efficiency. Moreover, the training speed of algorithms is greatly impacted by the scaling in MAS. We selected Humanoid(17x1) scenario, which involves 17 agents, for evaluating our methods under large-scale agents. In this scenario, the training time of diffusion-based methods significantly exceeds that of CPIG, demonstrating the scalability of our method in MARL.
@@
Due to the substantial reduction in sampling steps, the consistency model is inherently less expressive than the diffusion model [39]. To evaluate this, we conduct a performance comparison between CPIG and diffusion-based methods in three sparse scenarios. In comparison to CPIG-DM and CPIG-DPM, CPIG exhibits a performance decrease in MPE scenarios, yet outperforms it in the HalfCheetah(2x3) and Humanoid (17x1) environments. This phenomenon could potentially be attributed to that the consistency policy can effectively transmit action gradients in one step, thereby alleviating the challenges of policy training via Q-function in diffusion-based methods [37]. Moreover, the slight performance trade-off in CPIG is considered acceptable, given its significant improvements in training and inference efficiency.
@@ V-D Qualitative Analysis and Visualization
Figure 5: Example results obtained by CPIG and baselines on exploration visitation. (a) shows a Reacher4 task, with four targets in different colors. (b)-(d) shows the state visitation of CPIG, HASAC and MAPPO. (e) counts the targets covered by policies during exploration.
To showcase the exploration capabilities of our method, consider a didactic example illustrated in Fig. 5(a), which demonstrates different MARL algorithms applied to the sparse Reacher4 task [26] featuring four targets. The agents consist of two-arm robots tasked with moving the robot’s end effector close to one of the four targets. The depth of color represents the magnitude of the state visitation probability. In this scenario, we configure sparse rewards, meaning that the agents receive rewards only upon making contact with all four targets. Besides, the agents need to cooperate in exploration to reach the target, and the dispersion of targets often leads to agents adopting one single policy, underscoring the necessity for multimodal policies and cooperative exploration. In Fig. 5(b), our method learns a multimodal joint policy, enabling the effector to reach four different targets relatively uniformly, without converging to a unimodal policy. Conversely, in Fig. 5(c) and (d), baselines primarily focus on visiting certain targets, thereby limiting exploration in the complex environment. Additionally, we quantify the targets covered by the policy during exploration in Fig. 5(e). The curves indicate that our method achieves a state visitation rate of 83%, compared to 62% for HASAC, reflecting a 20% improvement. These results demonstrate that CPIG can facilitate a more efficient exploration of the four different targets in the scenario, highlighting the importance of multimodal policies and cooperative exploration capabilities among agents.
Figure 6: Trajectory visualization in HalfCheetah (2x3). Different colors represent the corresponding intention between agents.
@@
Figure 7: Visualization of the observation embeddings. We choose the last $30k$ steps for visualization. The colors denote different intentions.
In order to gain a better understanding of the intention learner within CPIG, we generate visualizations of intention guidance and observation embeddings. In Fig. 6, we visualize the true trajectories generated by the intention-guided consistency policy in the HalfCheetah(2x3) scenario. The trajectories are color-coded to represent the corresponding intention generated during execution. Across stages A, B, and C, although the observations of the two agents differ, they converge to a common intention (the same color). This indicates that the intention effectively reflects the global state, enabling multiple agents to coordinate their actions. By leveraging this shared intention, multiple agents can cooperate more effectively, thereby enhancing their exploration of the environment. Additionally, we present the visualizations obtained by applying the PCA technique to the observation embeddings. The goal is to show whether the intention learner can deduce distinguishable intention from the local observations, aiding in the guidance of policy during the exploration process. As shown in Fig. 7, the visualizations demonstrate that the intention learner can generate informative discrete representations by identifying potential correlations within individual observations of agents.
@@ V-E Ablation Study
In this subsection, we conduct ablation studies in the Navigation scenario to investigate the impact of three main components in CPIG: (1) replacing the consistency policy with a deterministic policy, similar to MADDPG [8], while directly incorporating intention guidance and observations directly as the inputs (CPIG-w/o-CP), (2) removing the intention learner module, forcing agents to rely solely on their own observations for exploration during training (unguided exploration) and for achieving objectives during execution (CPIG-w/o-IG), and (3) eliminating the self-reference mechanism, which relaxes the constraints on consistency policies (CPIG-w/o-SR). The experimental results, as shown in Fig. 8, demonstrate that CPIG-w/o-CP quickly converges to a locally optimal policy, highlighting the importance of the consistency policy in enabling effective exploration and preventing premature convergence to suboptimal solutions. Additionally, CPIG achieves superior performance compared to CPIG-w/o-IG, which lacks intention guidance, as evidenced by its faster convergence rate and higher final performance. Similarly, when using only intention guidance, CPIG-w/o-SR exhibits a lack of exploitation of past successes, reducing sample efficiency and leading to performance degradation. Overall, these findings underscore the importance of all three components—consistency policy, intention learner, and self-reference mechanism—in driving the improved performance of CPIG.
@@ VI Conclusions and Future Work
In this paper, we propose CPIG, a novel multi-agent exploration method that incorporates an intention-guided consistency policy to enable cooperative exploration and a self-reference mechanism to constrain generated actions. To the best of our knowledge, this study represents the first effort to leverage consistency policies in the context of MARL. Empirical results from six environments demonstrate that CPIG achieves performance comparable to baseline algorithms in dense reward settings while in sparse reward settings, which pose more challenges for exploration, our algorithm outperforms baselines by 20%.
This work is the first to showcase the benefits of intention-guided consistency policies in the realm of multi-agent exploration. However, the present work is confined to exploration within single-task scenarios. In the future, we intend to explore how to better harness the multimodal nature of consistency policies to devise a faster and more efficient algorithm for multi-task MARL.